OpenStudy (anonymous):

(x-5) (x+25) (5x+1)=0

OpenStudy (lalaly):

x=5 x=-25 x=-1/5

OpenStudy (anonymous):

x=-25,-1/5,5

OpenStudy (anonymous):

x=5 x=-25 x=-1/5

OpenStudy (anonymous):

solve

OpenStudy (anonymous):

x-5=0 (transpose -5) x=5

OpenStudy (anonymous):

x+25=0 (transpose 25) x=-25

OpenStudy (anonymous):

5x+1=0 5x=-1 x=-1/5

OpenStudy (anonymous):

x-5=0, (add 5 to both sides): x-5+5=0+5, x=5 x+25 = 0, (subtract 25 from both sides): x+25-25=0-25, x=-25 5x+1=0, (subtract 1from both sides): 5x+1-1=0-1, 5x=-1, (divide both sides by 5), 5x/5=-1/5, x=-1/5